3. Measurements shows that 500 muons fall on the top of Mt. Washington, altitude 2000m. Muons travel at 99% the speed of light or 0.99 x 3.0 x 108 m/s.

a. Calculate the time in microseconds it would take muons to travel from the top of Mt. Washington to its base.

b. Use your calculation and the muon clock graph to find how many muons should reach the bottom of Mt. Washington.


1
Expert's answer
2022-03-17T09:36:13-0400

"t=\frac sv=6.73~\mu s,"

"N=N_0(1-2^{-\frac tT})=475."